A number is 24 more than another. If their sum is 38, what are the numbers?
Hubert is three years older than Jovie. The sum of their ages is 51. Jovie's age is...

Answer by Alan3354(69443)

A number is 24 more than another. If their sum is 38, what are the numbers?
x + y = 38
x - y = 24
------------- Add
===============================
Hubert is three years older than Jovie. The sum of their ages is 51. Jovie's age is...
H = J + 3
H + J = 51
Sub for H in the 2nd eqn:
J+3 + J = 51
etc